OpenStudy (ineedhelpnow):

how do i write the point-slope for and the slope-intercept form of the equation of the line passing through (-1,60 with the slope of -4?

OpenStudy (anonymous):

y = mx + c m=-4 y =-4x +c (-1,6) 6 - 4=c =2 y=-4x+2

OpenStudy (anonymous):

y-y^1=m(x-x^1) y-60=-4(x-(-1) y-60=-4x-4 y=-4x+56

OpenStudy (anonymous):

I believe the points were (-1,6) as he must have had problem with shift key when pressing 0

OpenStudy (anonymous):

Oh thank you for clarifying, so it would be... y-y^1=m(x-x^1) y-6=-4(x-(-1) y-6=-4x-4 y=-4x+2